# FUSION IS ALL YOU NEED : FACE FUSION FOR CUSTOMIZED IDENTITY-PRESERVING IMAGE SYNTHESIS

### v1

@@ 3.3 MULTIPLE REFERENCE
Overall, we believe that a single image clearly showing the identity instance is sufficient for identitypreserving generation. However, there are cases where using multiple references can yield better results, especially when the face in the adapter is occluded or different features are displayed across different reference images. Our method can easily accommodate the use of multiple references. For identities $f1,\dots,fn$ with corresponding hidden states $h1,\dots,hn$, the keys and values can be obtained using Equation 2. Equation 3 is then modified as follows:
$K`=[K_{c};K_{f1};\dots;K_{fn}],V`=[V_{c};V_{f1};\dots;V_{fn}]$ (6)
@@
This modification allows for the integration of any number of identities without the need for further training or fine-tuning. We showcase sample results in Figure 3, demonstrating the effectiveness of using three identities for improved generation. The same principle can be applied to generate an image with an identity that is a mixture of distinct identities, a process commonly known as face morphing. More details on this application can be found in the appendix. By extending our method to handle multiple references, we enhance the flexibility and robustness of the identity-preserving generation process. This capability allows the model to leverage information from multiple sources, leading to better handling of occlusions and variations in facial features across different images.
@@ 3.4 MULTIPLE IDENTITIES
Handling multiple identities in image generation remains an open challenge. In this work, we propose a novel solution, that combines our previously mentioned method to address this issue. The goal is to generate an image containing (n) distinct identities ${f1,\dots,fn}$ based on input references and a given prompt. Given masks ${m1,\dots,mn}$ where mi indicates The location of The face for identity i, We scale each mask and feed to each cross-attention layer In The UNet, The scaling issue done based one the scale of the hidden state input To the attention and for some layer with scales si We have masks ${m1si,\dots,mnsi}$. For each of the face identities We can get get keys And values And combined them similar to Equation 6, if We applied attention like this we will obtain faces that are mix off all identities features (i.e. face morphing), we instead use a customized attention mask M, where The mask can is defined as:
$M_{q,k}=\begin{cases}-\inf&\text{if }isFaceKey(k)\times identity(k)=i\wedge\exists m_{j}:m_{j_{q}}=true\wedge j\neq i\\ 1&\text{otherwise }\end{cases}$ (7)
where q indicates the query and k is the key. This approach ensures that all identities are present in the generated image. We demonstrate the effectiveness of our method with graphical results showing example input identities in Figure 4. These results highlight the success of our approach in generating images conditioned on multiple face identities.
Figure 3: The overall pipeline of our proposed method. We first concatenate the target face image with the image to be generated, We modify the cross attention layers to include new keys any values layers their output is concatenated with the output of the original ones from text prompt.
@@ 4 Experiments
Figure 4: Multiple face identities image generation using our proposed method, first row represents the three input distinct identities, following row contains sample results for different prompts, depthmap based ControlNet Zhang et al. (2023) was used for the pose.
Figure 5: Comparison with other SOTA methods, with prompt used as ”astronaut in a garden” and the first row represents the input target faces identities, the following rows are results from IP-faceAdapter Ye et al. (2023), IP-face-Adapter-plus Ye et al. (2023) and ours respectively. Identities are selected from ffhq datasetKarras et al. (2019)
@@ 4.1 EXPERIMENT SETTINGS
In our work, we utilize Stable Diffusion Rombach et al. (2022) as the base model, specifically the Realistic Vision v4.0 checkpoint1 , we train a subset from the LAION-face dataset Zheng et al. (2022) consisting of approximately 80000 text-image pairs. Our training and all the experiments are done on 2 Tesla V100 GPUs (30 GB) with batch size of 1 per GPU. For the other models compared in our work, we use InstantID Wang et al. (2024) which uses SDXL model (offering an advantage over our method), as for IPA-FaceId, we use IPA-FaceID-plusv2 Ye et al. (2023) and both using Realistic Vision v4.0 checkpoint as the base model, for judging the CLIP score, we generate listing of 100 identity conditioned on different prompts using LLMs. We provide two versions of our method, the second version is trained for more timestep and larger dataset.
@@ 4.2 Evaluation Metrics
Previous research on identity-preserving imagine generation tasks Wang et al. (2024); Ye et al. (2023); Papantoniou et al. (2024) lacks a defined metric for assessing identity preservation. We employ cosine distance to assess the ID-conditioning capability of various approaches in our study. Face embeddings are acquired by feeding generated images into pre-trained face recognition models. We compare the face embedding of the reference image with those from the created image; a reduced distance signifies superior ID-consistent generating performance. Additionally, we calculate the CLIP score using the CLIP modelRadford et al. (2021) to assess content alignment in face image generation across several scenarios with varied text prompts. The higher CLIP score indicates high  [^1] https://huggingface.co/SG161222 text-image alignment. For the evaluation of image quality, we employ SSIM and PSNR to compare the generated images with the reference images. In order to measure the face fidelity, we focus on calculate those metrics based on the face region only by cropping the faces from generated images.
Figure 6: Comparison with other SOTA methods, the first column are the input identities along with input styles, From left to right, we show the results of IPA Ye et al. (2023), IPA-FaceID-SDXL, IPAFaceID-Plus, InstantId and our’s, it’s noticed that previous methods unlike our’s can suffer from lack of Face fidelity or style degradation.
@@ 4.3 QUALITATIVE RESULTS
In this section, we mainly evaluate our proposed method for visual comparison in terms of ID preservation. As showns in Figure 5, we generate images based on multiple identity face images with the same text prompt ”astronaut in a garden”. IPA-FaceID tend to generate face images with low ID preservation and it sometimes fails to maintain the gender of the reference face image. On the other hand, IPA-FaceID-Plus has better performance on maintaining the gender, age of input identity, however, it still loses some original facial attributes such as facial emotion and hairstyles. In the contrast, our method generates face images have high level of ID preservation (preserving facial expression, gender, age and detailed feature such as hairstyles and skincolor). Specifically, in order to investigate the accurate facial expression generation, we generate images with different text prompts like ”Anger”, ”Cry”, ”Laugh” to control the face emotion of generation. As depicted in Figure 1, InstantID generates images with less emotion alignment with associated text prompt while IPA-FaceID-Plus can generate intend emotion. However, IPA-FaceID-Plus has less control of ID preservation with exaggerated face expression in random styles (noticeable changes in facial structures and diverse backgrounds). Our method provides the strong control of facial expression based on corresponding text prompt. Especially in the case of ”Cry” prompt, even though the reference image already with ‘smile’ or ‘laugh’ facial expression, our method still can address the importance of text prompt to the final image generation. Our method produces ID-consistent face images while has more natural-looking face expression representations without suppressing the contribution of text prompt to the generated images. In Figure 6, we aim to compare the ID preservation in different image styles generation. Same as previous work Wang et al. (2024), we use a set of text prompts to indicate various image styles like ”Line art”, ”Faded Film” and ”Watercolor”. Our method is able to generate comparable results with other SOTA methods while maintaining high degree of face and style alignment together.
@@ 4.4 QUANTITATIVE RESULTS
Besides visual comparisons, we also numerically evaluate our method based on CLIP score, SSIM, PSNR and cosine distance using three models namely VGG-Face, Facenet512 and ArcFace. We utilize a subset of FFHQ Karras et al. (2019) dataset high quality face images as the input identities in these experiments.
Table 1 shows the numerical results of cosine distance on different methods. Firstly, in terms of the face alignment, it’s clear that our methods achieve significantly higher alignment than the other methods on all face recognition models. The faces in our generated image maintain high level of face features from original reference identity and can be easily used for accessing different face recognition systems. In Table 2, we judge The face quality And prompt alignment offering different methods, IPA-FaceId has the highest CLIP score, however it’s ID similarity is very low compared together ours and other methods, Ours v1 And IPA-FaceID-Plus maintains almost similar CLIP score with IPA-FaceID-Plus being slightly better, Ours v2 is worse In terms of CLIP but it achieves significantly higher identifier similarity. We exclude InstantId Wang et al. (2024) from CLip Score due to the fact that it requires pose conditioning And the base model for the public checkpoint issue SDXL based, which will make The comparison not plausible. We also compare The face quality using PSNR and SSIM metrics, our method achieves benefits result on both further demonstrating that our method is capable of generating high quality realistic faces with high ID fidelity.
@@ 5 LIMITATIONS AND FUTURE WORK
We have observed that fine facial features tend to degrade when the face occupies a small portion of the overall image. This issue primarily stems from the limitations of Stable Diffusion Rombach et al. (2022). Additionally, when our method is used in conjunction with ControlNet Zhang et al. (2023), it results in reduced identity fidelity, which can begin resolved by finetuning. In our future work, we aim to address these challenges by enhancing the ability to maintain fine details and improving control in scenarios involving the generation of multiple identities.
@@ 6 CONCLUSION
In this paper, we present a novel method for ID-preserving image generation that excels in maintaining high identity fidelity while producing realistic facial images. Our approach demonstrates robust generalization across various facial views and expressions. Additionally, we propose innovative solutions for handling multiple references and multi-identity generation, effectively bridging encoder domain gaps and leveraging multi-scale guidance for enhanced performance. Our method achieves new state-of-the-art results in preserving identity similarity, as evidenced by our extensive experimental evaluations. Furthermore, it maintains good alignment with given prompts, setting a new benchmark in the field of identity-preserving image generation.
Our proposed methods is capable of high prompt alignment with multi-views Face deduction from one sample input, while preserving the target identity with high fidelity 2, we can clearly see that other methods can fail in identity preservation on certain difficult reference inputs and loses prompt alignment, generating same face pose And landmarks is also another issue, we can see that our proposed pipeline and method of training overcome these issue as shown in 5.

### v2

@@ 3.3 Multiple References
Overall, we believe that a single image clearly showing the identity instance is sufficient for identity-preserving generation. However, there are cases where using multiple references can yield better results, especially when the face in the reference input is occluded or different features are displayed across different reference images. Our method can easily accommodate the use of multiple references. For identities $f_{1},\dots,f_{n}$ with corresponding hidden states $h_{1},\dots,h_{n}$, the keys and values can be obtained using Equation 2 and Equation 3 is then modified as follows:
$K^{{}^{\prime}}=[K_{c};K_{f_{1}};\dots;K_{f_{n}}],V^{{}^{\prime}}=[V_{c};V_{f_{1}};\dots;V_{f_{n}}]$ (6)
@@
This modification allows for the integration of any number of identities without the need for further training or fine-tuning. We showcase sample results in Figure 3, demonstrating the effectiveness of using three identities for improved generation. The same principle can be applied to generate an image with an identity that is a mixture of distinct identities, a process commonly known as face morphing (Fu & Damer, 2022), It is important to note that the order of the identities matters, as the attention mechanism in SD (Rombach et al., 2022) does not account for this.
By extending our method to handle multiple references, we enhance the flexibility and robustness of the identity-preserving generation process. This capability allows the model to leverage information from multiple sources, leading to better handling of occlusions and variations in facial features across different images.
@@ 3.4 Multiple Identities
Figure 4: Multiple face identities image generation using our proposed method, first row represents the three input distinct identities, following row contains sample results for different prompts, depthmap based ControlNet (Zhang et al., 2023) was used for the pose.
Handling multiple identities in image generation remains an open challenge. In this work, we propose a solution, that combines our previously mentioned method to address this issue. The goal is to generate an image containing (n) distinct identities $f_{1},\dots,f_{n}$ based on input references and a given prompt. Given masks $M_{1},\dots,M_{n}$ where $M_{i}$ indicates the location mask of the face for identity i, we scale each mask and feed to each cross-attention layer in the UNet, the scaling is done based on the scale of the hidden state input to the attention and for some layer with scales s the reshaped and scaled masks are ${M_{1}}^{(s)},\dots,{M_{n}}^{(s)}$, for each of the face identities, we can get keys and values that are combined following Equation 6, if we applied attention like this we will obtain faces that are mix of all identities features (i.e. face morphing), we instead use a customized attention mask, denoted as ${\bm{A}}$, where the mask is defined as:
${\bm{A}}_{q,k}=\begin{cases}-\inf&\text{if }isFaceKey(k)\wedge identity(k)=i\wedge\exists{M_{j}}^{(s)}:{M_{j_{q}}}^{(s)}=true\wedge j\neq i\\ 1&\text{otherwise }\end{cases}$ (7)
where q indicates the query and k is the key, this restricts the guidance for each face reference to a single, specific face, it also ensures that all identities are present in the generated image. We demonstrate the effectiveness of our method with graphical results using distinct input identities in Figure 4. These results highlight the success of our approach in generating images conditioned on multiple face identities.
@@ 4 Experiments
Figure 5: Comparison with other SOTA methods, with prompt used as “astronaut in a garden” and the first row represents the input target faces identities, the following rows are results from IP-faceAdapter (Ye et al., 2023), IP-face-Adapter-plus (Ye et al., 2023) and ours respectively. Identities are selected from ffhq dataset (Karras et al., 2019).
@@ 4.1 Experiment Settings
In our work, we utilize Stable Diffusion (Rombach et al., 2022) as the base model, specifically the Realistic Vision v4.0 checkpoint [^1] https://huggingface.co/SG161222/Realistic_Vision_V4.0_noVAE. We train on a subset of the LAION-face dataset (Zheng et al., 2022), consisting of approximately 80,000 text-image pairs. All training and experiments are conducted on two Tesla V100 GPUs (32 GB), with a batch size of 1 per GPU. For comparison with other models, we use InstantID (Wang et al., 2024), which employs the SDXL model (providing an advantage over our method). For IPA-FaceId, we use IPA-FaceID-plusv2 (Ye et al., 2023), both utilizing the Realistic Vision v4.0 checkpoint as the base model. To evaluate the CLIP score, we generate a list of 100 identities conditioned on different prompts using LLMs. We provide two versions of our method, with the second version trained for more timesteps and on a larger dataset.
@@ 4.2 Evaluation Metrics
Previous research on identity-preserving imagine generation tasks (Wang et al., 2024; Ye et al., 2023) lacks a defined metric for assessing identity preservation. We employ cosine distance to assess the ID-conditioning capability of various approaches in our study. Face embeddings are acquired by feeding generated images into pre-trained face recognition models. We compare the face embedding of the reference image with those from the created image; a reduced distance signifies superior ID-consistent generating performance. Additionally, we calculate the CLIP score using the CLIP model (Radford et al., 2021) to assess content alignment in face image generation across several scenarios with varied text prompts, the higher CLIP score indicates high text-image alignment. For the evaluation of image face quality, we employ SSIM and PSNR to compare the generated images with the reference images. In order to measure the face fidelity, we focus on calculating those metrics based on the face region only by cropping the faces from the generated images.
Figure 6: Visual comparison with other SOTA methods in different generation styles. The first column are the input identities along with input styles, from left to right, we show the results of IPA (Ye et al., 2023), IPA-FaceID, IPA-FaceID-SDXL, IPA-FaceID-Plus, InstantID (Wang et al., 2024) and ours. It’s noticed that previous methods suffer from the lack of face fidelity or style degradation. Images are selected from InstantID (Wang et al., 2024).
@@ 4.3 Qualitative Results
In this section, we evaluate our proposed method for visual comparison in terms of identity (ID) preservation. As shown in Figure 5, we generate images based on multiple identity face images with the same text prompt “astronaut in a garden”. IPA-FaceID tends to generate face images with low ID preservation and sometimes fails to maintain the gender and age of the reference face image. On the other hand, IPA-FaceID-Plus performs better in maintaining the gender and age of the input identity; however, it still loses some original facial attributes, such as skin tone and hairstyles.
In contrast, our method generates face images with a high level of ID preservation, preserving facial expressions, gender, age, and detailed features such as hairstyles and skin color. Specifically, to investigate accurate facial expression generation, we generate images using different text prompt styles like “Anger”, “Cry”, and “Laugh” to control the facial emotion in the generated images. As depicted in Figure 6, InstantID generates images with less emotional alignment to the associated text prompts, while IPA-FaceID-Plus can generate the intended emotion. However, IPA-FaceID-Plus has less control over ID preservation, producing exaggerated facial expressions in random styles. Our method provides strong control of facial expressions based on the corresponding text prompt, especially in the case of the “Cry” prompt, even if the reference image already has a ‘smile’ or ‘laugh’ facial expression, our method still emphasizes the importance of the text prompt in the final image generation. Our method produces ID-consistent face images while delivering more natural-looking facial expression representations, without suppressing the contribution of the text prompt to the generated images.
In Figure 6, we aim to compare the ID preservation in different image styles generation. Same as previous work InstantID (Wang et al., 2024), we use a set of text prompts to indicate various prompt styles like “Line art”, “Faded Film” and “Watercolor”. Our method is able to generate comparable results with other SOTA methods while maintaining high degree of face and style alignment together.
Table 1: Quantative evaluations of ID preservation.
Cosine distance $\downarrow$
Method VGG-faces Facenet512 ArcFace
IPA-FaceID 0.60 0.27 0.55
IPA-FaceID-Plus 0.36 0.25 0.39
InstantID 0.42 0.27 0.38
$Ours_{v1}$ 0.33 0.22 0.34
$Ours_{v2}$ 0.25 0.18 0.28
@@ 4.4 Quantitative Results
Besides visual comparisons, we also numerically evaluate our method based on CLIP score, SSIM, PSNR and cosine distance using deepface (Serengil & Ozpinar, 2024; 2020) utilizing three models namely VGG-Face, Facenet512 and ArcFace. We utilize a subset of FFHQ (Karras et al., 2019) dataset high quality face images as the input identities in these experiments.
Table 1 shows the numerical results of cosine distance on different methods. Firstly, in terms of the face alignment, it’s clear that our methods achieve significantly higher alignment than the other methods on all face recognition models. The faces in our generated image maintain high level of face features from original reference identity and can be easily used for deceiving different face recognition systems.
In Table 2, we judge the face quality and prompt alignment offering different methods, IPA-FaceId has the highest CLIP score, however its’ ID similarity is very low compared to ours and other methods, ours v1 and IPA-FaceID-Plus maintains almost similar CLIP score with IPA-FaceID-Plus being slightly better, ours v2 is worse in terms of CLIP, but it achieves significantly higher identity similarity. We exclude InstantID (Wang et al., 2024) from CLIP score due to the fact that it requires pose conditioning and the base model for the public checkpoint is SDXL based, which will make the comparison not plausible. We also compare the face quality using PSNR and SSIM metrics, our method achieves best result on both, further demonstrating that our method is capable of generating high quality realistic faces with high ID fidelity.
@@
We have observed that fine facial features tend to degrade when the face occupies a small portion of the overall image. This issue primarily stems from the limitations of Stable Diffusion (Rombach et al., 2022). Additionally, when our method is used in conjunction with ControlNet (Zhang et al., 2023), it results in slight reduction in identity fidelity, which can be resolved by light fine-tuning or reducing the conditioning strength. In our future work, we aim to address these challenges by enhancing the ability to maintain fine details and improving control in scenarios involving the generation of multiple identities.
@@ 6 Conclusion
In this paper, we present a novel method for ID-preserving image generation that excels in maintaining high identity fidelity while producing realistic facial images. Our approach demonstrates robust generalization across various facial views and expressions. Additionally, we propose innovative solutions for handling multiple references and multi-identity generation, effectively bridging encoder domain gaps and leveraging multi-reference guidance for enhanced performance. Our method achieves new state-of-the-art results in preserving identity similarity, as evidenced by our extensive experimental evaluations. Furthermore, it maintains good alignment with given prompts, setting a new benchmark in the field of identity-preserving image generation.
Table 2: Quantative evaluations of face quality and prompt alignment.
Method CLIP Score $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
IPA-FaceID 25.23 27.94 0.36
IPA-FaceID-Plus 23.76 28.04 0.44
InstantID - 27.93 0.33
$Ours_{v1}$ 23.28 28.25 0.41
$Ours_{v2}$ 22.51 28.40 0.47